find three consecutive integers whose sum is 9 greater than twice the largest number
Answer by oberobic(2304)

Define the numbers as:
x
x+1
x+2
.
x +(x+1) +(x+2) = 2(x+2) + 9
.
3x + 3 = 2x + 4 + 9
.
x = 10
x+1 = 11
x+2 = 12
.
Check the answer.
2*12 + 9 = 24+9 = 33
10+11+12 = 33
Correct.
.
Answer: The three integers are 10, 11, and 12.
.
Done.